# code[Code] DistMLIP: A Distributed Inference Platform for Machine Learning Interatomic Potentials

### v1

@@ 4 Results
In this section, we benchmark DistMLIP with the 4 MLIPs loaded with their public pretrained checkpoints: MACE-MP-0b-small-3.8M, CHGNet-2.7M, TensorNet-0.8M, and eSEN-3.2M. The details of the models and checkpoints can be found in appendix C.
@@ 4.1 Maximum Capacity
We also benchmark the maximum capacity and corresponding inference time against the SevenNet model, the results can be found in Appendix G. After matching the total number of parameters to SevenNet (800k), we find that MACE, TensorNet, and CHGNet can achieve up to 10x higher maximum capacity and 4x faster inference speed when incorporated with DistMLIP compared to the distributed inference of SevenNet.
@@ 4.2 Strong and Weak Scaling
Weak scaling tests, where the total size of the system increases proportionally with the number of GPUs (such that each GPU performs computation on the same number of atoms), are found in Figure 2(c). All times are normalized to be represented as multiples of the 1 GPU time. The per GPU atom count for MACE-3.8M, TensorNet-0.8M, CHGNet-2.7M, and eSEN-3.2M are 34.6k, 19.9k, 9.9k, and 1.4k, respectively. In the case of eSEN-3.2M, weak scaling moves away from ideal scaling due to the high constant overhead associated with initial feature calculation. For CHGNet-2.7M, the computation required for the construction of the three-body graph scales with where is the number of atoms within the three-body cutoff, leading to suboptimal weak scaling when the simulation cell size increases.
@@ 4.3 Interaction Range
The results in Fig. 3 (a) show that DistMLIP only has a linear relation between parallelized inference time vs. interaction range. This is due to the additional computation cost from each increased message passing layer. Conversely, in conventional space partitioning, the volume of the simulation cell, and therefore the number of ghost atoms, grows cubically with the interaction range. This highlights the parallelization efficiency and zero calculation redundancy in graph partitioning.
@@ 4.4 Scaling Model Size
In Fig. 3 (b) and (c), we fix the number of message passing layers and vary the feature embedding sizes in the MLIP, therefore measuring the relation between parallelized inference speed/capacity and model parameter size. The result shows that by decreasing the model parameter size, a significant increase in simulation speed and maximum capacity can be achieved. The result suggests an estimated performance gain when distributed inference can be combined with smaller model sizes through MLIP model distillation (Amin et al., 2025).
@@ 4.5 Real World Simulations
We also show the performance of real distributed simulations on a variety of solid-state and biomolecular systems, utilizing 1, 4, and 8 GPUs. The results are found in Table 1. We report the microsceonds/atom-timestep of each model-system pair as well as the number of simulated atoms in the system. The simulated systems can be found in Figure 4. L-MACE-3.8M refers to multi-GPU inference of MACE using LAMMPS spatial partitioning, while the other 4 models are distributed with DistMLIP. Note that L-MACE-3.8M uses a compiled model with custom equivariant CUDA kernels, while DistMLIP MACE-3.8M only runs the pure-PyTorch implementation. Custom equivariant CUDA kernels have been shown to accelerate MACE inference time by up to 7.2x on large models (Geiger et al., 2024). No other model beyond MACE is reported due to the lack of LAMMPS multi-GPU inference support.
@@
Our result shows that most of the current foundational potentials, even at significant model sizes of a few million parameters, are capable of simulating near-million-atom scale systems when parallelized with DistMLIP on 8 GPUs. Moreover, we noticed that the inference time, when normalized by the number of atoms, is significantly decreased when any MLIP is being parallelized. This observation suggests chemically rare events can be cheaply simulated using a larger cell for a shorter simulation time, rather than a smaller cell for a longer simulation time, which is the standard way of current atomistic simulations. As estimated from the benchmark result in table 1, nanosecond near-million-atoms simulations can be achieved at the order of 10 days with standard MLIPs and DistMLIP on a few GPUs. The max capacity of the 12Å-MACE is tripled when coupled with DistMLIP compared to LAMMPS space partitioning. For a longer-range model like CHGNet, the capacity increase and inference speed-up will be more significant as memory usage in spatial partitioning scales cubically with interaction range.
@@ 5 Conclusion
Scaling the quantum-chemical simulation accuracy to realistic application scales remains a critical challenge even with recent developments of ML interatomic potentials. To address this challenge, we present DistMLIP, a distributed MLIP inference platform based on efficient graph-level partitioning. Compared to the traditional parallelization method, space-partitioning through LAMMPS, DistMLIP serves as an easy and versatile distributed inference platform that supports long-range MLIPs. DistMLIP provides infrastructures for constructing and distributing atom and bond graphs, allowing the distribution of GNN-based MLIPs that are otherwise infeasible to parallelize.
We benchmarked the parallelized inference of 4 popular MLIPs: MACE, TensorNet, CHGNet and eSEN. Our result shows that efficient and plug-and-play parallelization can be achieved when combining DistMLIP with existing interatomic potentials. By distributing the MLIP simulation on 8 NVIDIA-A100 GPUs, our result shows that nanosecond, near-million-atom scale simulations can be accomplished at the scale of 10 physical days with state-of-the-art MLIPs. We believe this effort to enable large-scale simulation would accelerate chemical, materials, and biological discovery.
Acknowledgments and Disclosure of Funding
@@
Appendix A Distributing Three-body Graphs
Algorithm 2 depicts the method in which to distribute three-body graphs (also known as bond graphs) as well as calculating the necessary information to perform data transfer between various partitions at each convolution of the three-body graph.
@@ Appendix C MLIP Versions in Benchmark
The table below shows the checkpoint versions of the MLIPs tested. Note that the eSEN model we benchmarked is not coming from the public pretrained checkpoints of 30.2M parameters, which is too big for efficient parallelized simulation. Instead, we initialized a 3.2M eSEN in accordance with the eSEN-MPTrj-3.2M configuration found in Fu et al. [2025].
@@
| CHGNet | matgl-MatPES-PBE-2025.2.10 | 2.7M | 45Å | [Deng et al., 2023] |
| MACE | MACE-MP-0b-small | 3.8M | 12Å | [Batatia et al., 2023] |
| TensorNet | matgl-MatPES-PBE-v2025.1 | 0.8M | 10Å | [Ko et al., 2025] |
| eSEN | eSEN-MPTrj-3.2M | 3.2M | 12Å | [Fu et al., 2025] |
@@ Appendix D Single GPU Benchmarking Details
Because DistMLIP parallelizes neighbor list construction as well as underlying threebody graph creation, utilizing 2 DistMLIP partitions oftentimes leads to faster total inference time and less total memory consumption compared to a baseline implementation (this is especially the case with CHGNet). Therefore, to maintain a fair comparison, all single-GPU results reported in any benchmark utilize 2 DistMLIP partitions performing operations on the same GPU. Therefore, only 1 GPU is utilized, but the same fast graph creation algorithms and implementation are shared. For all benchmarking tasks, 128 threads were used for neighbor list construction and graph creation.
@@ Appendix E Inference Time Breakdown
Neighbor list construction could take a substantial amount of inference time when the simulated system is large. In order to address this issue, we parallelized neighbor list construction in DistMLIP through multi-threading, so that graph creation time is substantially decreased compared to the single-thread neighbor list construction in Pymatgen [Ong et al., 2013]. The resulting inference time breakdown is shown in Table 3.
| Stage | MACE | TensorNet | CHGNet |
| Graph Creation | 0.2480 | 0.1262 | 0.3386 |
| Feature Calculation | 0.0437 | 0.1297 | 0.1018 |
| Forward Pass | 0.1701 | 0.1782 | 1.8216 |
| Backward Pass | 0.0288 | 0.5471 | 1.4074 |
@@ Appendix H Alignment of Single-device and Multi-device Predictions
DistMLIP’s atom graph and bond graph distribution algorithms are exact in principle. However, numerical differences arise when performing computation in a distributed manner compared to on a single GPU. This is a result of non-determinism occurring during matrix multiplications and other operations on different GPUs. Therefore, the exact same model and weights running single GPU inference on different GPUs within the same node will also yield slightly different results. In Fig. 7, we plot the energy/atom error in meV/atom units for MACE-3.8M, TensorNet-0.8M, and CHGNet-2.7M. The result shows that the numerical error from different GPUs is far below chemical accuracy.
Appendix I Parallelizing a Model in DistMLIP
DistMLIP is designed to be both high-performant as well as easily usable. Parallelizing new MLIPs using DistMLIP is a straightforward process that can be done purely in Python. A few key points of DistMLIP are outlined in I.1.

### v2

@@ 4 Results
In this section, we benchmark DistMLIP with the 4 FPs loaded with their public pretrained checkpoints: MACE-MP-0b-small-3.8M, CHGNet-2.7M, TensorNet-0.8M, and eSEN-3.2M. The details of the models and checkpoints can be found in appendix C.
@@ 4.1 Maximum Capacity
We also benchmark the maximum capacity and corresponding inference time against the SevenNet model. The results can be found in Appendix G. After matching the total number of parameters to SevenNet (800k), we find that MACE, TensorNet, and CHGNet can achieve up to 10x higher maximum capacity and 4x faster inference speed when incorporated with DistMLIP compared to the distributed inference of SevenNet.
@@ 4.2 Strong and Weak Scaling
Weak scaling tests, where the total size of the system increases proportionally with the number of GPUs (such that each GPU performs computation on the same number of atoms), are found in Figure 2(c). All times are normalized to be represented as multiples of the 1 GPU time. The per GPU atom count for MACE-3.8M, TensorNet-0.8M, CHGNet-2.7M, and eSEN-3.2M are 34.6k, 19.9k, 9.9k, and 1.4k, respectively. In the case of eSEN-3.2M, weak scaling moves away from ideal scaling due to the high constant overhead associated with initial feature calculation. For CHGNet-2.7M, the computation required for the construction of the three-body graph scales with where is the number of atoms within the three-body cutoff, leading to suboptimal weak scaling when the simulation cell size increases. In Table 3 of Appendix K.1, we show that the simple vertical partitioning rule used in DistMLIP and specified in Algorithm 3 is up to 8x faster compared to standard graph partitioning baselines.
@@ 4.3 Interaction Range
The results in Fig. 3 (a) show that DistMLIP only has a linear relation between parallelized inference time vs. interaction range. This is due to the additional computation cost from each increased message passing layer. Conversely, in conventional spatial partitioning, the volume of the simulation cell, and therefore the number of ghost atoms, grows cubically with the interaction range. This highlights the parallelization efficiency and zero calculation redundancy in graph partitioning.
@@ 4.4 Scaling Model Size
In Fig. 3 (b) and (c), we fix the number of message passing layers and vary the feature embedding sizes in the MLIP, therefore measuring the relation between parallelized inference speed/capacity and model parameter size. The result shows that by decreasing the model parameter size, a significant increase in simulation speed and maximum capacity can be achieved. The result suggests an estimated performance gain when distributed inference can be combined with smaller model sizes through MLIP model distillation (Amin2025Distill).
@@ 4.5 Real World Simulations
We also show the performance of real distributed simulations on a variety of solid-state and biomolecular systems, utilizing 1, 4, and 8 GPUs. The results are found in Table 1. We report the microsceonds/atom-timestep of each model-system pair as well as the number of simulated atoms in the system. The simulated systems can be found in Figure 4. In Table 1, L-MACE-3.8M refers to multi-GPU inference of MACE using LAMMPS spatial partitioning, while the other 4 models are distributed with DistMLIP.
@@
Our result shows that DistMLIP provides tripled maximum simulation sizes compared to LAMMPS spatial partitioning within the MACE-3.8M model. Note that L-MACE-3.8M uses a compiled model with custom equivariant CUDA kernels, while DistMLIP MACE-3.8M only runs the pure-PyTorch implementation. Custom equivariant CUDA kernels were shown to accelerate MACE inference time by up to 7.2x on large models (geiger2024cuequivariance). Nevertheless, we observed similar simulation speed between the standard model on DistMLIP and the compiled model on LAMMPS, which supports the efficiency of DistMLIP graph-partitioning. No other model beyond MACE is reported due to the lack of LAMMPS multi-GPU inference support. For other FPs, which typically have longer interaction ranges compared to 12Å in MACE-3.8M, the capacity increase and inference speed-up should be much more significant as the efficiency of spatial partitioning degrades rapidly with increased cutoffs.
In Table 1, we highlight that most FPs with a few million parameters are capable of simulating near-million-atom scale systems when parallelized with only 8 GPUs. Moreover, we noticed that the inference time, when normalized by the number of atoms, is significantly decreased when any MLIP is being parallelized. This observation suggests chemically rare events can be cheaply simulated using a larger cell for a shorter simulation time, rather than a smaller cell for a longer simulation time, which has been the standard simulation procedure due to the inability to efficiently perform large simulations. As estimated from the benchmark result in Table 1, nanosecond near-million-atoms simulations can now be achieved at the order of 10 days with standard FPs and DistMLIP on a few GPUs.
@@ 5 Conclusion
Scaling quantum-chemical simulations to the size of realistic applications remains a critical challenge, even with recent developments of MLIPs and FPs. To address this challenge, we present DistMLIP, a distributed MLIP inference platform based on efficient graph-level partitioning. Compared to the conventional spatial partitioning through LAMMPS, DistMLIP serves as an easy and versatile distributed inference platform that supports long-range MLIPs. DistMLIP provides infrastructures for constructing and distributing atom and bond graphs, allowing the distribution of GNN-based MLIPs that are otherwise infeasible to parallelize.
We benchmarked the parallelized inference of 4 popular MLIPs: MACE, TensorNet, CHGNet and eSEN. Our result shows that efficient and plug-and-play parallelization can be achieved when combining DistMLIP with existing interatomic potentials. By distributing the MLIP simulation on 8 NVIDIA-A100 GPUs, our result shows that nanosecond, near-million-atom scale simulations can be accomplished at the scale of 10 physical days with state-of-the-art FPs. We believe this effort to enable large-scale simulation would accelerate chemical, materials, and biological discovery.
Acknowledgements and Disclosure of Funding
@@
Appendix A Distributing Bond Graphs
Algorithm 2 depicts the method to distribute three-body graphs (bond graphs), as well as calculating the necessary information to perform data transfer between various partitions at each convolution of the three-body graph.
@@ Appendix C MLIP Versions in Benchmark
The table below shows the checkpoint versions of the MLIPs tested. The CHGNet model is taken from recent release of MatGL library Ko2025Matgl. The eSEN model in our benchmark is not taken from the public pretrained checkpoints of 30.2M parameters, which is too big for efficient parallelized simulation. Instead, we initialized a 3.2M eSEN in accordance with the eSEN-MPTrj-3.2M configuration found in esen.
@@
| CHGNet | matgl-MatPES-PBE-2025.2.10 | 2.7M | 45Å | (deng2023chgnet) |
| MACE | MACE-MP-0b-small | 3.8M | 12Å | (Batatia_2023) |
| TensorNet | matgl-MatPES-PBE-v2025.1 | 0.8M | 10Å | (Ko2025Matgl) |
| eSEN | eSEN-MPTrj-3.2M | 3.2M | 12Å | (esen) |
@@ Appendix D Single GPU Benchmarking Details
Because DistMLIP parallelizes neighbor list construction as well as underlying threebody graph creation, utilizing only 2 DistMLIP partitions can already lead to faster total inference time and less total memory consumption compared to a baseline implementation without DistMLIP (this is especially the case with CHGNet). Therefore, to maintain a fair comparison, all single-GPU results reported in any benchmark utilize 2 DistMLIP partitions performing operations on the same GPU. Therefore, only 1 GPU is utilized, but the same fast graph creation algorithms and implementation are shared. For all benchmarking tasks, 128 threads were used for neighbor list construction and graph creation.
@@ Appendix E Inference Time Breakdown
Neighbor list construction could take a substantial amount of inference time when the simulated system is large. In order to address this issue, we parallelized neighbor list construction in DistMLIP through multi-threading, so that graph creation time is substantially decreased compared to the single-thread neighbor list construction in Pymatgen (Ong2013Pymatgen). Furthermore, we show the breakdown of overall inference time using DistMLIP when fixing the total atomic system size in Figure 5. We also include the breakdown of overall inference time when fixing the total number of atoms per GPU while scaling the total number of GPUs in Figure 6. The atomic system used was a crystalline SiO2 supercell expanded in a cubic fashion.
@@ Appendix H Alignment of Single-device and Multi-device Predictions
DistMLIP’s atom graph and bond graph distribution algorithms are exact in principle. However, numerical differences arise when performing computation in a distributed manner compared to on a single GPU. This is a result of non-determinism occurring during matrix multiplications and other operations on different GPUs (fatahalian2004understanding). Therefore, the exact same model and weights running single GPU inference on different GPUs within the same node will also yield slightly different results. In Fig. 9, we plot the energy/atom error in meV/atom units for MACE-3.8M, TensorNet-0.8M, and CHGNet-2.7M. The result shows that the numerical error from different GPUs is far below chemical accuracy.
Appendix I Usage
Appendix J Parallelizing a Model in DistMLIP
DistMLIP is designed to be both high-performant as well as easily usable. Parallelizing new MLIPs using DistMLIP is a straightforward process that can be done purely in Python. A few key points of DistMLIP are outlined in J.1. The primary data structure, the Distributed object, inputs atom positions, periodic boundary conditions, and (optionally) cell lattice, and constructs the graph partitions and associated metadata. The data structure can partition, aggregate, and perform data transfer for node features, edge features, and node features within the threebody graph. These simple distributed primitives are implemented in high performance C and makes distributing a new MLIP very straightforward to implement but still high performing.
J.1 Model Parallelization Example
Appendix K Graph Partitioning
Graph partitioning algorithms find applications in solving PDEs via domain decomposition, solving sparse linear systems of equations, circuit partitioning and layout, VLSI design, social network analysis, clustering algorithms, and image segmentation (pothen1997graphpartapplications; stanton2012streaming; bader2013graphclustering; tolliver2006imagegraph; peng2013imagegraph). One common use case in graph partitioning is to create mutually exclusive spanning sets of nodes that contain the minimum number of edges that cross from one partition to another. This use case can be tackled using sparse, symmetric matrix reordering methods such as the Reverse Cuthill-Mckee (RCM) algorithm, which permutes a sparse matrix to minimize its bandwidth (i.e. reordering rows and columns such that non-zero values are closer to the diagonal) (cuthillmckee; azad2017reverse). The reordered matrix can then be partitioned along the columns in order to calculate graph node partitions. METIS is a graph partitioning algorithm that utilizes a graph coarsening phase, an initial partitioning sequence over the coarsened graph, and an uncoarsening and partition refinement stage (metis). However, applications depending on algorithms such as RCM or METIS typically don’t have latency requirements during the graph partitioning stage. In atomistic simulation, the underlying graphs are recalculated and repartitioned at each time step. Therefore, even small latency increases during the graph creation and partitioning stage get compounded into significant increases in overall simulation time. In our own experiments, we find that RCM and METIS could increase inference time for million-atom graphs by several seconds per timestep.
K.1 Benchmarking Partition Strategies
We compare DistMLIP’s vertical wall partitioning strategy with other common graph partitioning algorithms. In Table 3, we replace DistMLIP’s current vertical wall partitioning strategy with the Reverse Cuthill-McKee (RCMK) and METIS algorithms while holding the other components of Algorithm 1 and Algorithm 3 constant. Neither RCMK nor METIS supports threebody bond graph creation. RCMK and METIS both utilize the graph’s topology in order to partition the graph such that the number of crossing edges between partitions is minimized. DistMLIP’s current partitioning strategy, on the other hand, doesn’t perform graph traversals but rather uses atomic positions as a heuristic in order to partition the graph. In Table 3, we also include the LAMMPS spatial partitioning results for comparison. As a result, we perform all benchmarks with the MACE-3.8M model.
| Method | # GPUs | / (atomstep) # of atoms (in thousands) | ||||
| Li3PO4 | H2O | GaN | MOF | 2w49 | ||
| METIS | 4 GPUs | |||||
| 8 GPUs | ||||||
| RCMK | 4 GPUs | |||||
| 8 GPUs | ||||||
| Vert. wall | 4 GPUs | |||||
| 8 GPUs | ||||||
Appendix L Molecular Dynamics Simulation Stability
To validate the numerical robustness and long-term stability of DistMLIP simulations, we performed a 2 nanosecond TensorNet MD simulation of a Li-ion cathode material containing 0.1 million atoms on 8 A100 GPUs. Fig. 10 shows the energy of the material as a function of time. Throughout the simulation, the system maintained structural stability. Except for the expected thermal fluctuation, no severe energy oscillation is observed during the entire simulation. The initial decrease in energy is not a numerical artifact; rather, it is attributed to the energy equilibration of the simulated material, representing a crucial physical process successfully captured by the DistMLIP simulation.
Appendix M Large system training using DistMLIP
Machine learning interatomic potentials are trained on quantum mechanical calculations such as density functional theory or coupled cluster techniques. Due to the computational complexity of these techniques, however, calculating the energy and forces of atomic systems with greater than several hundred atoms is computationally intractable. As a result, other than instances in which large batch sizes are used in training, DistMLIP’s primary use case would be for large scale inference. However, for technical completeness, we perform training on a GPCR protein, 6P9X from the protein data bank (berman2000protein) consisting of 8.1k atoms. Energy and forces were calculated using a Lennard-Jones potential. With a batch size of 16, we achieve 2.1 seconds per training step on 8 A100-80 GB GPUs using a 0.4M parameter CHGNet model – demonstrating DistMLIP’s ability in large-scale, large-batch training.